OpenStudy (anonymous):

verify that the numbers 1+ sqrt 5 and 1- sqrt5 both satisfy the equation x^2-2x-4=0

OpenStudy (amoodarya):

\[s=x1+x2 =(1-\sqrt{5})+(1+\sqrt{5})=2\\p=x1*x2 =(1-\sqrt{5})*(1+\sqrt{5})=-4\\x^2-sx+p\]

OpenStudy (anonymous):

so it doesn't satisfy it?

OpenStudy (anonymous):

put these values in the equation

OpenStudy (anonymous):

or wait i have a better method make the quadratic using the roots given and compare it with the quadratic equation given

OpenStudy (anonymous):

\[x^2-(1+\sqrt5+1-\sqrt5)x+(1-\sqrt5)(1+\sqrt 5)=0\]

OpenStudy (anonymous):

\[x^2-2x-4=0\]

OpenStudy (anonymous):

okay makes sense thanks